OpenStudy (liizzyliizz):

Equation of tangent line to the graph f(x)= (x-2)(x^2-3x-1) at the point (-1,-9) is the equation y+9=9(x+1) or am i missing a step

OpenStudy (amistre64):

whats your f' ?

OpenStudy (amistre64):

f(x)= (x-2)(x^2-3x-1) f'(x)= (x-2)'(x^2-3x-1) + (x-2)(x^2-3x-1)' = (1)(x^2-3x-1) + (x-2)(2x-3) f'(-1) = ((-1)^2-3(-1)-1) + ((-1)-2)(2(-1)-3) = (1+3-1) + (-3)(-5) = 3 + 15 = 18 or did i miss the math?

OpenStudy (amistre64):

the rest looks good; change the slope tho

OpenStudy (liizzyliizz):

oh ok that must of been my math then. i see where i went wrong i just realized i missed one term when using the product rule. but thank you !

OpenStudy (amistre64):

;) youre welcome